OpenStudy (kwonbon):

Will give medal! Please help~ Find the remaining trigonometric ratios. tan(a)=2,0 < a < pi/2 sin(a)= cot(a)= csc(a)= cos(a)= sec(a)=

OpenStudy (loser66):

Where are you stuck?

jhonyy9 (jhonyy9):

use tan = sin /cos

jhonyy9 (jhonyy9):

do you know what mean csc and sec ? and csc = ? and sec = ?

OpenStudy (kwonbon):

@jhonyy9 @Loser66 I know what each individual ratio is (csc=1/sin, etc.), but I just don't know how to go about solving the problem :(

jhonyy9 (jhonyy9):

i ve wrote above that tan = sin / cos use this

OpenStudy (kwonbon):

@jhonyy9 , so... I still don't understand what tan would be. Would it be 2?

OpenStudy (kwonbon):

@jhonyy9 if it is 2, then sin/cos = 2/1... but the answer said that sin does not equal 2

jhonyy9 (jhonyy9):

sin ---- = 2 cos this not is equivalent with sin 2 --- = ---- and from this result than sin = 2 and cos = 1 cos 1 hope you understand it now

jhonyy9 (jhonyy9):

and hope you know that cot = 1/tan

OpenStudy (kwonbon):

But the answer said that sin doesn't equal 2 @jhonyy9

OpenStudy (kwonbon):

I'm still confused lol

jhonyy9 (jhonyy9):

yes this i right sin not equal 2 bc. sin/cos = 2

jhonyy9 (jhonyy9):

sin 2 --- = ---- cross multiply and will get sin = 2cos cos 1

OpenStudy (kwonbon):

so... sin=2cos... But you don't know what cos is? @jhonyy9

OpenStudy (kwonbon):

nvm I figured it out lol but your answer didn't help at all

jhonyy9 (jhonyy9):

just you need using your knowledge from trigonometry - yes ?

jhonyy9 (jhonyy9):

for example cot = 1/tan so cot = 1/2 yes ?

OpenStudy (kwonbon):

mhm

jhonyy9 (jhonyy9):

@OtherWorldly something about this please ?

OpenStudy (otherworldly):

i don't know trigonometry sorry :(

OpenStudy (otherworldly):

haven't gotten to that point yet

OpenStudy (otherworldly):

@jackthegreatest pretty sure u would know this

OpenStudy (jackthegreatest):

yep i would